alones (alones):

I NEDE help I still can't figure out this :What is the slope of the line that is perpendicular to the line whose equation is 2x + y = 4.

alones (alones):

I actually got 1/2 :/

OpenStudy (dezandalf42714):

what are the choices.

OpenStudy (anonymous):

y= -2x +4 then the perpendicular is y=1/2x +4

alones (alones):

-2 -1/2 1/2 2

OpenStudy (anonymous):

you get y= -2x+4 from the original equation you had, then change the inverse of the original slope which is y=1/2x +4

alones (alones):

o_O i get it ,i didn't even think of that solution thx :)

OpenStudy (anonymous):

YEah @DeonAkki is right :P